OpenStudy (anonymous):

Differentiate: g(x) = (1+4x)^5 * (3 + x - x^2)^8

OpenStudy (anonymous):

I got as far as 8 (4x+1)^5 (-2x+1) (-x^2+x+3)^7 + 20 (4x+1)^4 (-x^2+x+3)^8, but the textbook's answer is 4(1+4x)^4 (3+x-x^2)^7 (17+9x-21x^2) .....

OpenStudy (bahrom7893):

g'(x)=20(1+4x)^4*(3x+x-x^2)^8+8(1-2x)(1+4x)^5(3+x-x^2)^7

OpenStudy (bahrom7893):

wait that 3x is supposed to be 3. Typing long expressions is conusing.

OpenStudy (bahrom7893):

g'(x)=20(1+4x)^4*(3+x-x^2)^8+8(1-2x)(1+4x)^5(3+x-x^2)^7

OpenStudy (anonymous):

yeah but i have to get that to the textbook answer

OpenStudy (bahrom7893):

g'(x)=4(1+4x)^4(3+x-x^2)^7[5(3+x-x^2)+2(1-2x)(1+4x)]

OpenStudy (bahrom7893):

g'(x)=4(1+4x)^4(3+x-x^2)^7[15+5x-5x^2+2(1+4x-2x-8x^2)]

OpenStudy (anonymous):

but how

OpenStudy (bahrom7893):

pull out a 4, (1+4x)^4 and (3+x-x^2)^7, those are all common terms

OpenStudy (anonymous):

you mean 4, (1+4x) and (3+x-x^2)?

OpenStudy (bahrom7893):

No, to the 7th power.. one is to the 7th and one is to the 8th so just pull out the whole thing

OpenStudy (anonymous):

riight wait so divide both sides by 4 (3+x-x^2)^7 (1+4x)^4 ??